# Self-Supervised Skeleton Action Representation Learning: A Benchmark and Beyond

### v1

@@ Abstract
Self-supervised learning (SSL), which aims to learn meaningful prior representations from unlabeled data, has been proven effective for label-efficient skeleton-based action understanding. Different from the image domain, skeleton data possesses sparser spatial structures and diverse representation forms, with the absence of background clues and the additional temporal dimension. This presents the new challenges for the pretext task design of spatial-temporal motion representation learning. Recently, many endeavors have been made for skeleton-based SSL and remarkable progress has been achieved. However, a systematic and thorough review is still lacking. In this paper, we conduct, for the first time, a comprehensive survey on self-supervised skeleton-based action representation learning, where various literature is organized according to their pre-training pretext task methodologies. Following the taxonomy of context-based, generative learning, and contrastive learning approaches, we make a thorough review and benchmark of existing works and shed light on the future possible directions. Our investigation demonstrates that most SSL works rely on the single paradigm, learning representations of a single level, and are evaluated on the action recognition task solely, which leaves the generalization power of skeleton SSL models under-explored. To this end, a novel and effective SSL method for skeleton is further proposed, which integrates multiple pretext tasks to jointly learn versatile representations of different granularity, substantially boosting the generalization capacity for different downstream tasks. Extensive experiments under three large-scale datasets demonstrate that the proposed method achieves the superior generalization performance on various downstream tasks, including recognition, retrieval, detection, and few-shot learning.
Index Terms:
@@ 1 Introduction
In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [4, 5, 6]. However, these supervised methods heavily rely on huge amounts of labeled data, which requires time-consuming and expensive data annotation work, limiting the wide applications in the real world. As a remedy to this problem, self-supervised learning (SSL) attracts much attention and has been proven successful for representation learning. It exploits supervisory signals from unlabeled data, learning meaningful prior features and boosting generalization capacity of model for downstream tasks. Motivated by recent success in the image domain, great interest has arisen in adopting SSL for skeleton. However, it is not trivial to transfer these approaches into the skeleton data directly, which are with a more compact spatial structure, additional temporal dimension, and the absence of the background clues. To this end, researchers have made valuable exploration for skeleton-based SSL.
@@
  • We present a comprehensive benchmark of existing skeleton SSL works as well as a summary of the popular datasets and downstream tasks for SSL evaluation. Meanwhile, we provide insightful analysis from the perspective of model backbones and pre-training paradigms, and discuss the future possible directions.
@@
  • We perform rigorous quantitative experiments to study the generalization efficacy of current skeleton-based SSL methods across five downstream tasks, including recognition, retrieval, detection, and few-shot learning, on both uncorrupted and corrupted skeletons. Under the proposed SSL method, the promising results are achieved for versatile action representation learning.
@@
2 Review on Self-Supervised Skeleton-Based Action Representation Learning
@@
2.1 Skeleton Data as Human Representation
@@ 2.2.1 Context-Based Methods
1) View Invariance. Due to the variance of the observation viewpoint, the skeleton estimation can suffer from occlusion and noise. Therefore, learning view-invariant representations should be beneficial for action recognition task, and has been widely studied in the supervised skeleton-based action recognition task [30]. Li et al[31] proposed a view classification pretext task for unsupervised action representation learning. Specifically, as shown in Fig. 3 (a), skeleton sequences of different views are fed into the encoder, subsequent to which a view classifier predicts the view labels. To learn the view-invariant features, a Gradient Reversal Layer [32] is added to reverse the optimization direction of the encoder. Resort to this, the encoder can learn the features insensitive to view in an adversarial manner. Likewise, Paoletti et al[33] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
@@
3) Motion Prior. The motion dynamics of skeleton joints contain rich semantic information beneficial for action understanding. Therefore, researchers propose to leverage the inherent motion prior knowledge to generate pseudo labels for pre-training as shown in Fig. 3 (c). Cheng et al[40] designed a movement direction prediction task. The model is constrained to estimate the direction of the instantaneous joint velocity, i.e., whether the targeted joints are moving in a positive direction. Based on this, Kim et al[41] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning. The key idea is to predict the motion direction and magnitude of the central joint and other joints during a random interval. On the other hand, Yang et al[42] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation. The model is pre-trained by a regression task to predict this prior knowledge.
@@ 2.2.2 Generative Learning Methods
When $\mathcal{T}_{in}$ and $\mathcal{T}_{tar}$ are both the identity function, the model is constrained to reconstruct the input data simply, known as AutoEncoder (AE). It naturally creates an information bottleneck to achieve dimension reduction, mapping data from input space onto low-dimensional feature space. Due to its simpleness, AE is used to learn prior representations in earlier skeleton-based works [14, 43]. However, since the skeleton data is still redundant in spatial and temporal dimensions, the encoder tends to memorize the input data at a low level instead of modeling the high-level semantic knowledge. Then, denoising AE (DAE) is studied for alleviating this shortcut [44, 45]. For example, Nie et al[45] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features. On the other hand, inspired by the remarkable success of masked language/image modeling (MLM, MIM), which can be regarded as a variant of DAE, researchers have aroused a surge of interest in exploring masked skeleton modeling (MSM) for human action representation learning. The earlier works are mainly based on recurrent neural network (RNN) or graph convolutional network (GCN) to predict the masked skeletons. However, lacking good scalability, these models only show mediocre performance improvement. Recently, inspired by the success of Vision Transformer (ViT) as masked autoencoders (MAE) [46], Transformer-based models have been explored for the masked skeleton modeling [47, 48]. However, due to the relative redundancy in skeleton and the lack of a large-scale dataset as ImageNet, directly applying Transformer for masked skeleton reconstruction can suffer from the over-fitting problem. To this end, researchers have made endeavors on two crucial designs, masking strategy and reconstruction target, which are discussed as follows.
@@
Instead of directly reconstructing the input, Mao et al[48] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames. The work [54] formulates the skeleton as an unordered 3D point cloud and maps the 3D data onto color space. This mapping function is artificially defined according to the spatial-temporal relationship of skeleton joints. Therefore, the model can learn the spatial relation and temporal dependency by reconstructing the color of the skeleton cloud.
2.2.3 Contrastive-Learning-Based Methods
@@
For skeleton contrastive representation learning, most existing methods are based on MoCo v2 [49]. It adopts an asynchronously momentum-updated key encoder and an online query encoder, along with a memory queue to store a large number of consistent negative samples. Specifically, different augmentations are applied to the skeleton $x$ to generate the positive pair $(x_{q},x_{k})$, while the negative sample features $m_{i}$ are stored in a memory queue. The model is constrained to retrieve the positives among the negative samples, optimizing the following InfoNCE objective:
$\mathcal{L}_{Info}(z_{q},z_{k})=-\log\frac{\exp(z_{q}\cdot z_{k}/\tau)}{\exp(z_{q}\cdot z_{k}/\tau)+\sum_{i=1}{\exp(z_{q}\cdot m_{i}/\tau)}},$ (2)
@@
1) Augmentation Design and Learning Strategy. Data augmentation exposes novel motion patterns and generates diverse positive views, which have been found crucial to the success of contrastive learning [59, 60]. Different from images, data augmentation for skeleton is relatively less developed. Therefore, the earlier works mainly focus on exploring practical spatial-temporal augmentations for skeleton contrastive learning [61, 62, 63]. For example, Rao et al[61] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research. Notably, the optimal augmentations for different backbones are usually different. For example, the Joint Mask is found detrimental for GCNs [12], while beneficial for GRU model [62], which implies that GCNs are more sensitive to the spatial corruption. Generally, this difference arises on account of different modeling mechanisms and model capacities, which further increases the difficulty of the augmentation design and selection.
Inspired by the success of mixing-based augmentations for images, Chen et al[64] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences. To learn from this mixed skeleton, SkeleMixCLR obtains the corresponding part level and the whole-body level features and pursues the local-global invariance. Based on this, SkeAttnCLR [65] further designs an attention mechanism to perform local contrastive learning on salient and non-salient features. Due to the generated novel input views and regularization effect on the feature space, mixing augmentation often leads to consistent improvement for representations.
@@
where $z_{weak}$ and $z_{strong}$ are the corresponding embeddings of weakly and strongly augmented views. Instead of the one-hot target, this objective utilizes the similarity distribution as the soft target to guide consistency learning, which can be viewed as a self-distillation process. Further, Zhang et al[12] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework. It performs a decoupled progressive augmentation invariance learning by optimizing the consistency only between the augmented samples with adjacent strength, where the weakly augmented branch serves as the mimic target of the strongly augmented branch. These works show that the weakly augmented views can effectively guide the learning of the corresponding strongly augmented samples, leading to more stable representation learning and improvement.
In addition, Lin et al[67] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy. By mining the actionlet in an unsupervised manner, the semantic-reserving augmentations are employed for actionlet regions, while the noise perturbations for non-actionlet regions, avoiding the over-distortion problem. On the other hand, instead of augmenting at the input level, HaLP [68] proposes a latent positive hallucinating method by exploring the latent space around the corresponding prototype.
@@ 2.3 Summary and Discussion
3 The Proposed Method: Prompted Contrast with Masked Motion Modeling
@@ 4 Benchmark Skeleton SSL Methods
TABLE I: Summary of popular skeleton datasets. Marker denotes the dataset is collected by Marker-MoCap methods.
Dataset’Year Instance Class. Sub. View Joint Marker
CMU [82]’03 2235 45 144 - 31
SBU [83]’12 300 8 7 - 15
NW-UCLA [84]’14 1,494 10 10 3 20
UWA3D [85]’15 1,075 30 10 5 15
SYSU [86]’15 480 12 40 1 20
NTU 60 [26]’16 56,880 60 40 80 25
PKUMMD [28]’17 21,545 51 66 3 25
Kinetics [87]’17 $\sim$260k 400 - - 18
TSU [88]’19 16,115 31 18 - 17
NTU 120 [27]’19 114,480 120 106 155 25
UAV [89]’21 22,476 155 119 17
BABEL-60 [90]’21 45473 60 - - 25
BABEL-120 [90]’21 48,978 120 - - 25
@@ 4.1 Datasets
We first summarize the popular datasets for skeleton action understanding in Table I. Among them, the following datasets are widely used in skeleton-based SSL evaluation.
@@
4) Northwestern-UCLA (NW-UCLA) [84] contains 1,494 action samples, captured by three Kinect v1 cameras. Samples of 10 action categories are included, performed by 10 subjects. Following the recommended protocol, the training data are from the first two cameras while the testing data are from the other one.
@@ 4.3 Downstream Tasks
1) Action Recognition is the most common evaluation task for skeleton-based SSL works. Typically, two evaluation protocols are widely adopted. The first is the linear evaluation protocol, where a linear layer is added with the pre-trained model fixed. The other is fine-tuning protocol where the whole model is trained including the subsequent linear layer. Top-k accuracy metric is adopted in this task.
@@ 4.4 Implementations Details
For the existing methods, we give priority to the reported results in their original paper. While for the reproduced results, we strictly keep the fairness of the comparison and implementation. Detailed settings of the reproduction experiments can be found in the supplementary material.
For our proposed method PCM3++, we adopt the three-layer Bi-GRU with the hidden dimension of $d=1024$ as the encoder backbone and follow the experiment settings including the data pre-processing and pre-training of our baseline method [71], which are consistent with other GRU-based works [10, 62] for comparison fairness. During the pre-training stage, the model is trained for 450 epochs with additional 150 epochs for post refinement. The batch size is 128. We set learning rate as 0.02, which is reduced to 0.002 at the 350th epoch. The SGD optimizer is adopted with a momentum of 0.9. $\tau$, $\tau_{q}$, and $\tau_{k}$ are set to 0.07, 0.1, and 0.05, respectively. $\lambda_{m}$ and $\lambda_{kd}$ are 40.0 and 1.0. We conduct all experiments on an NVIDIA RTX 4090 GPU.
@@ 4.5 Benchmark on Action Recognition Task
4.5.1 Comparison across Different Pretext Tasks
As shown in Fig. 5, context-based and generative learning pretext tasks are mainly studied in the earlier works first. However, these methods usually fail to achieve a satisfactory performance under linear evaluation due to the modeling of too many low-level features and the lack of effective design. Recently, MAMP [48] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol, which means that Transformer-based generative learning is promising.
@@ 4.5.1 Comparison across Different Pretext Tasks
4.5.2 Comparison across Different Model Backbones
Different model backbones often employ different pre-training tasks. For example, Transformer often consumes extensive computational resources, making it difficult for contrastive learning, which often requires encoding multiple data views. In contrast, it naturally fits the MSM schema which reduces computational overhead by masking tokens with large ratios. On the other hand, GRU and GCN models are relatively more efficient in training, often adopting the contrastive learning pretext task. It is also noted that the GRU models are often with larger feature dimensions than GCN models due to the smaller induced computational graph and less GPU memory occupancy.

### v2

# Self-Supervised Skeleton-Based Action Representation Learning: A Benchmark and Beyond

@@ Abstract
Self-supervised learning (SSL), which aims to learn meaningful prior representations from unlabeled data, has been proven effective for skeleton-based action understanding. Different from the image domain, skeleton data possesses sparser spatial structures and diverse representation forms, with the absence of background clues and the additional temporal dimension, presenting new challenges for spatial-temporal motion pretext task design. Recently, many endeavors have been made for skeleton-based SSL, achieving remarkable progress. However, a systematic and thorough review is still lacking. In this paper, we conduct, for the first time, a comprehensive survey on self-supervised skeleton-based action representation learning. Following the taxonomy of context-based, generative learning, and contrastive learning approaches, we make a thorough review and benchmark of existing works and shed light on the future possible directions. Remarkably, our investigation demonstrates that most SSL works rely on the single paradigm, learning representations of a single level, and are evaluated on the action recognition task solely, which leaves the generalization power of skeleton SSL models under-explored. To this end, a novel and effective SSL method for skeleton is further proposed, which integrates versatile representation learning objectives of different granularity, substantially boosting the generalization capacity for multiple skeleton downstream tasks. Extensive experiments under three large-scale datasets demonstrate our method achieves superior generalization performance on various downstream tasks, including recognition, retrieval, detection, and few-shot learning.
keywords:
@@ 1 Introduction
In the early works, many endeavors have been put into the supervised skeleton-based human activity understanding, e.g., recognition and detection [86, 63, 11, 59]. However, these supervised methods heavily rely on huge amounts of labeled data, which requires time-consuming and expensive data annotation work, limiting the wide applications in the real world. As a remedy to this problem, self-supervised learning (SSL) attracts much attention and has been proven successful for representation learning. It exploits supervisory signals from unlabeled data, learning meaningful prior features and boosting generalization capacity of model for downstream tasks. Motivated by recent success in the image domain, great interest has arisen in adopting SSL for skeleton. However, it is not trivial to transfer these approaches into the skeleton data directly, which are with a more compact spatial structure, additional temporal dimension, and the absence of the background clues. To this end, researchers have made valuable exploration for skeleton-based SSL.
@@
  • We present a comprehensive benchmark of existing skeleton SSL works as well as a summary of the popular datasets and downstream tasks as well as the insightful analysis from the perspective of model backbones, pre-training paradigms, and future possible directions. Finally, we demonstrate promising performance of our proposed method on five downstream tasks for versatile action representation learning.
@@
2 Review on Skeleton-Based Action Representation SSL
@@
2.1 Human Skeleton Representation
@@ 2.2.1 Context-Based Methods
1) View Invariance. Due to the variance of the observation viewpoint, the skeleton estimation can suffer from occlusion and noise. Therefore, learning view-invariant representations should be beneficial for action recognition task, and has been widely studied in the supervised skeleton-based action recognition task [19]. Li et al. [37] proposed a view classification pretext task for unsupervised action representation learning. Specifically, as shown in Fig. 3 (a), skeleton sequences of different views are fed into the encoder, subsequent to which a view classifier predicts the view labels. To learn the view-invariant features, a Gradient Reversal Layer [18] is added to reverse the optimization direction of the encoder. Resort to this, the encoder can learn the features insensitive to view in an adversarial manner. Likewise, Paoletti et al. [55] adopted the rotation prediction, i.e., pitch angle, yaw angle, and roll angle, as well as a gradient reverse operation to achieve the viewpoint-invariance learning.
@@
3) Motion Prior. The motion dynamics of skeleton joints contain rich semantic information beneficial for action understanding. Therefore, researchers propose to leverage the inherent motion prior knowledge to generate pseudo labels for pre-training as shown in Fig. 3 (c). Cheng et al. [12] designed a movement direction prediction task. The model is constrained to estimate the direction of the instantaneous joint velocity, i.e., whether the targeted joints are moving in a positive direction. Based on this, Kim et al. [33] proposed a multi-interval pose displacement strategy to encourage both local and global attention learning. The key idea is to predict the motion direction and magnitude of the central joint and other joints during a random interval. On the other hand, Yang et al. [91] introduced more motion prior information as the pseudo label, i.e., intra-joint motion variance, inter-joint motion covariance, intra-frame joint angle, and inter-frame motion deviation. The model is pre-trained by a regression task to predict this prior knowledge.
@@ 2.2.2 Generative Learning Methods
When $\mathcal{T}_{in}$ and $\mathcal{T}_{tar}$ are both the identity function, the model is constrained to reconstruct the input data simply, known as AutoEncoder (AE). It naturally creates an information bottleneck to achieve dimension reduction, mapping data from input space onto low-dimensional feature space. Due to its simpleness, AE is used to learn prior representations in earlier skeleton-based works [68, 84]. However, since the skeleton data is still redundant in spatial and temporal dimensions, the encoder tends to memorize the input data at a low level instead of modeling the high-level semantic knowledge. Then, denoising AE (DAE) is studied for alleviating this shortcut [69, 54]. For example, Nie et al. [54] applied a series of view corruptions and constrained the model to reconstruct the clean data to disentangle the view and pose features. On the other hand, inspired by the remarkable success of masked language/image modeling (MLM, MIM), which can be regarded as a variant of DAE, researchers have aroused a surge of interest in exploring masked skeleton modeling (MSM) for human action representation learning. The earlier works are mainly based on recurrent neural network (RNN) or graph convolutional network (GCN) to predict the masked skeletons. However, lacking good scalability, these models only show mediocre performance improvement. Recently, inspired by the success of Vision Transformer (ViT) as masked autoencoders (MAE) [25], Transformer-based models have been explored for the masked skeleton modeling [9, 51]. However, due to the relative redundancy in skeleton and the lack of a large-scale dataset as ImageNet, directly applying Transformer for masked skeleton reconstruction can suffer from the over-fitting problem. To this end, researchers have made endeavors on two crucial designs, masking strategy and reconstruction target, which are discussed as follows.
@@
Instead of directly reconstructing the input, Mao et al. [51] proposed to reconstruct the motion in MSM, i.e., the difference of the corresponding joints between adjacent frames. The work [90] formulates the skeleton as an unordered 3D point cloud and maps the 3D data onto color space. This mapping function is artificially defined according to the spatial-temporal relationship of skeleton joints. Therefore, the model can learn the spatial relation and temporal dependency by reconstructing the color of the skeleton cloud.
2.2.3 Contrastive-Learning Methods
@@
For skeleton contrastive representation learning, most existing methods are based on MoCo v2 [26]. It adopts an asynchronously momentum-updated key encoder and an online query encoder, along with a memory queue to store a large number of consistent negative samples. Specifically, different augmentations are applied to the skeleton $x$ to generate the positive pair $(x_{q},x_{k})$, while the negative sample features $m_{i}$ are stored in a memory queue. The model is constrained to retrieve the positives among the negative samples, optimizing the following InfoNCE objective $\mathcal{L}_{Info}(z_{q},z_{k})$:
$\mathcal{L}=-\log\frac{\exp(z_{q}\cdot z_{k}/\tau)}{\exp(z_{q}\cdot z_{k}/\tau)+\sum_{i=1}{\exp(z_{q}\cdot m_{i}/\tau)}},$ (2)
@@
1) Augmentation Design and Learning Strategy. Data augmentation exposes novel motion patterns and generates diverse positive views, which have been found crucial to the success of contrastive learning [73, 24]. Different from images, data augmentation for skeleton is relatively less developed. Therefore, the earlier works mainly focus on exploring practical spatial-temporal augmentations for skeleton contrastive learning [60, 72, 20]. For example, Rao et al. [60] proposed a series of augmentations including Rotation, Shear, Reverse, Gaussian Noise, Gaussian Blur, Joint Mask and Channel Mask, some of which are used as the default basic augmentations in the future research. Notably, the optimal augmentations for different backbones are usually different. For example, the Joint Mask is found detrimental for GCNs [97], while beneficial for GRU model [72], which implies that GCNs are more sensitive to the spatial corruption. Generally, this difference arises on account of different modeling mechanisms and model capacities, which further increases the difficulty of the augmentation design and selection.
Inspired by the success of mixing-based augmentations for images, Chen et al. [94] proposed SkeleMixCLR equipped with SkeleMix augmentation which combines the topological information of different skeleton sequences. To learn from this mixed skeleton, SkeleMixCLR obtains the corresponding part level and the whole-body level features and pursues the local-global invariance. Based on this, SkeAttnCLR [29] further designs an attention mechanism to perform local contrastive learning on salient and non-salient features. Due to the generated novel input views and regularization effect on the feature space, mixing augmentation often leads to consistent improvement for representations.
@@
where $z_{weak}$ and $z_{strong}$ are the corresponding embeddings of weakly and strongly augmented views. Instead of the one-hot target, this objective utilizes the similarity distribution as the soft target to guide consistency learning, which can be viewed as a self-distillation process. Further, Zhang et al. [97] introduced more strong augmentations, e.g., randomly dropping the skeleton edges and joints, and proposed a hierarchical contrastive learning framework. It performs a decoupled progressive augmentation invariance learning by optimizing the consistency only between the augmented samples with adjacent strength, where the weakly augmented branch serves as the mimic target of the strongly augmented branch. These works show that the weakly augmented views can effectively guide the learning of the corresponding strongly augmented samples, leading to more stable representation learning and improvement.
In addition, Lin et al. [41] explicitly distinguished the static regions and motion regions, namely, actionlet, in human skeleton, and introduced a motion-aware augmentation strategy. By mining the actionlet in an unsupervised manner, the semantic-reserving augmentations are employed for actionlet regions, while the noise perturbations for non-actionlet regions, avoiding the over-distortion problem. On the other hand, instead of augmenting at the input level, HaLP [61] proposes a latent positive hallucinating method by exploring the latent space around the corresponding prototype.
@@ 2.3 Summary and Discussion
3 The Proposed Method
@@ 4 Benchmark Skeleton SSL Methods
Table 1: Summary of popular skeleton datasets. * denotes the collected by Marker-MoCap methods.
Dataset Instance Class Sub. View Joint
CMU* [13] 2235 45 144 - 31
SBU [92] 300 8 7 - 15
NW-UCLA [76] 1,494 10 10 3 20
UWA3D [58] 1,075 30 10 5 15
SYSU [28] 480 12 40 1 20
NTU 60 [62] 56,880 60 40 80 25
PKUMMD [43] 21,545 51 66 3 25
Kinetics [32] $\sim$260k 400 - - 18
TSU [14] 16,115 31 18 - 17
NTU 120 [44] 114,480 120 106 155 25
UAV [39] 22,476 155 119 17
BABEL* [56] 48,978 120 - - 25
@@ 4.1 Datasets
We first summarize the popular datasets for skeleton action understanding in Table 1. Among them, the following datasets are widely used in skeleton-based SSL evaluation.
@@ 4.3 Downstream Tasks
1) Action Recognition is the most common SSL evaluation task. Typically, two evaluation protocols are widely adopted. The first is the linear evaluation protocol, where a linear layer is added with the pre-trained model fixed. The other is fine-tuning protocol where the whole model is trained including the subsequent linear layer. Top-k accuracy metric is adopted in this task.
@@ 4.4 Implementations Details
For the proposed method, we follow the implementation and experiment settings of previous works [98, 52]. For the existing methods, we give priority to the reported results in their original paper. Meanwhile, we reproduce some results to keep the fairness of the comparison and implementation. Details can be found in Supp.
@@ 4.5 Benchmark on Action Recognition Task
1) Comparison across Different Pretext Tasks. As shown in Fig. 5, context-based and generative learning pretext tasks are mainly studied in the earlier works first. However, these methods usually fail to achieve a satisfactory performance under linear evaluation due to the modeling of too many low-level features and the lack of effective design. Recently, MAMP [51] adopts MSM with large mask ratios and motion-aware reconstruction targets, obtaining a highly discriminative semantic feature space for both linear evaluation and fully fine-tuning protocol, demonstrating promising results for Transformer-based generative learning.
@@
2) Comparison across Different Model Backbones. Different model backbones often employ different pre-training tasks. For example, Transformer often consumes extensive computational resources, making it difficult for contrastive learning, which often requires encoding multiple data views. In contrast, it naturally fits the MSM schema which reduces computational overhead by masking tokens with large ratios. On the other hand, GRU and GCN models are relatively more efficient in training, often adopting the contrastive learning pretext task. It is also noted that the GRU models are often with larger feature dimensions than GCN models due to the smaller induced computational graph and less GPU memory occupancy.